OpenStudy (anonymous):

f(x)=-x^2-3x+6 evaluate f(-2). Show all steps.

OpenStudy (anonymous):

Just substitute -2 for x

hero (hero):

f(x) = -x^2-3x + 6 f(-2) = -(-2)^2 - 3(-2) + 6 f(-2) = -4+6 + 6 f(-2) = -4 + 12 f(-2) = 8

hero (hero):

I knew he was going to run away when he got it wrong